a mass of 5 kg with velocity 4 m/s collides with a mass of 8 kg with velocity 3 m/s. it is divided into 2 masses of 3 kg and 10kg where 10 kg mass gined a velocity of 1 m/s. what will be new speed of obtained 3 kg mass ?

Grade:12th pass

1 Answers

Vikas TU
5 years ago
As no external force is acting upon the system, the momentum will be conserved.
pi = pf
mv = constant.
5*4 + 8*3 = 3*v + 10*1
20 + 24 = 3v + 10
3v = 44 – 10
3v = 34
v = 11.33 m/s